OpenStudy (anonymous):

change (13,6) to polar points

OpenStudy (anonymous):

also (-9,-6)

OpenStudy (zehanz):

You have to use \(r=\sqrt{x^2+y^2}\) and \( \theta = \arctan\frac{b}{a}\)

OpenStudy (anonymous):

i can find r but all my angles are nonreal

OpenStudy (zehanz):

Make that last one \( \theta=\arctan\frac{y}{x}\)

OpenStudy (anonymous):

k let me try

OpenStudy (zehanz):

My calculator gives 24.8 degrees for (13,6)

OpenStudy (anonymous):

i need radians though

OpenStudy (zehanz):

0.43

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about the second one

OpenStudy (zehanz):

THe calculator says 0.59 radians. That is wrong, because we're in the third quadrant here. So add pi radians to get there. Or: subtract pi, doesn't matter.

OpenStudy (anonymous):

so its .59+pi

OpenStudy (zehanz):

Yes, I would turn that into 3.73, because 0.59 isn't exact either.

OpenStudy (anonymous):

ok thanks

OpenStudy (zehanz):

YW!

OpenStudy (anonymous):

i still says i am wrong for (-9,-6) also (-4,6)